OpenStudy (anonymous):

help pleaseee What is the equation of the circle with center at (2, 4) and radius of 6?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

equation of circle with centre as (h,k) and radius r is (x-h)^2+(y-k)^2=r^2 just put (h,k)=(2,4) and r=6 and simplify you'll get the required equation..

OpenStudy (anonymous):

can you do it for me @abhi7

OpenStudy (anonymous):

(x-2)^2+(y-4)^2=6^2 use (a-b)^2=a^2-2ab+b^2 so x^2-4x+4+y^2-8y+16=36 or,x^2+y^2-4x-8y=16 which is the required equation